OpenStudy (anonymous):

one side of a rectangle is 1m less than the length of the diagonal. If the area of a rectangle is 60m squared, what are the dimensions ? answer plss

OpenStudy (gorv):

let the length = x metre so other side= x-1

OpenStudy (gorv):

area = x*(x-1) area is 60 it given so 60=x(x-1) 60=x^2-x subtract 60 from both side

OpenStudy (gorv):

x^2-x-60=0 now solve this quadratic